Hi,
My question is:
How many milliliters (mL) of water must be added to 10 milliliters of a 30%-sulfuric acid solution to make a 15%-sulfuric acid solution?

Thanks

Answer by lokesh mishra(5)

since in 10 milimeters 3 ml is sulphuric acid , let x ml be added ,(10 +x) 85/100 = (7+x) , 140 + 20x = 170 + 17x , 3x = 30 , x = 10 ml , of water sud b added